# Blockchain with proof of quantum work

### v1

@@ D.1 Blockchain Efficiency and Transaction Delay
We define blockchain efficiency as the expected number of blocks in the strongest path divided by the total number of blocks broadcast, in the limit of many blocks. For viable parameterizations this concentrates at a finite value as the number of block broadcasts becomes large. For purposes of Figs. 5 and 15 we estimate this value once after every block broadcast for blockchains of between $512$ and $2048$ broadcasts (mining events). We discard the values for the first 50% of the chain evolution and average the remaining result. The resulting value is shown as a mean and standard error with respect to either 5 or 10 independent chains.
@@
For figure 15 we measure 1 pair-delay per chain following each mining event on chains of length from 512 up to 2048. We determine the empirical distribution of $D_{pair}$, discarding statistics from the first half of each chain. We threshold the distribution to determine a transaction delay with 95% confidence, presenting the mean and variance with respect to at least 5 (at most 10) blockchains. Although pair-delays measured this way are correlated, for each data point we have at least thousands of effectively independent samples for each data point, so that a 95% quantile is established with reasonable confidence [49].
@@ D.2 Mining Acceleration for Blockchain Analysis
Each mining event requires $O(2^{\mathcal{N}_{\rm zeros}})$ experiments. For purposes of blockchain structural analysis we show methods whereby only $O(1)$ experiments are required per mining event. This allows us to present, with practical compute resources, results on stability for large $\mathcal{N}_{\rm zeros}$, limited primarily by the number of validation experiments rather than mining.
@@
We first make the approximation in all experiments of perfect network synchronization, discounting the rare event of two miners simultaneously completing a proof-of-work. This approximation is reasonable if the work is mining rate is small enough.
We employ a hashing method where the overall sign on any hyperplane is uniformly distributed in $\{-1,1\}$. For this reason the probability a nonce describes a broadcastable (valid) block in case (a) is $1/2^{-\mathcal{N}_{\rm zeros}}$; mining is successful if and only if the hyperplane signs $H_{\alpha}=\mathrm{sign}(W_{\alpha})$. We can therefore emulate properly distributed statistics of a fair mining event by first running an experiment to produce $W_{\alpha}$, and then resampling the signs on hyperplanes to allow mining.
@@
D.3 Resampling Witness Methodology
@@
For a training set of 5 experiments and $1024$ random projections, the distribution of bits (thence validation rates), and distribution of $W_{\alpha}$ (thence confidence in validation) can be established as shown in Figs. 12 and 13. We can fit the witness distributions by Gaussians, and thresholded bit outcomes by Bernouilli distributions. For purposes of accelerating analysis we can then resample these distributions to model a larger set of experiments without use of the QPU experiment each time.
We employ this procedure in Figs. 5 and 15 with results consistent with full blockchain experiments involving finite-number of miners, such as those in Fig. 9. We consider 5120 combinations of model (spin-glass realization) and hyperplanes, for each of the 4 experimental QPUs. Using 20 programmings, we determine the distribution (Gaussian or Bernouilli for witnesses and bits respectively). We can then use these parameterized distributions to generate reasonable witness (and thresholded witness) distributions. For each mining attempt, we uniformly randomly sample a combination of model and hyperplanes - and then for any mining or validation event we choose uniformly at random an QPU and sample from the corresponding distribution.
@@
Qualitative justification for this approach relies on the observation that a spin-glass ensemble operated in our experimental regime is self-averaging. The distributional properties of such witnesses fluctuates little between typical spin-glass disorder realizations. Variation in witnesses values between QPUs (and embeddings) appears well described by Gaussian distributions, and since hyperplanes are nearly orthogonal the correlations in their fluctuations are weak. However, we expect this method to underestimate fluctuations so that Figs. 5 and 15 may be biased toward higher efficiency (and reduced delay). Full simulations (every verification and mining event uses a QPU) are used for all other figures. The outcomes are in qualitative agreement with the resampling method, and reasonable quantitative agreement with respect to blockchain efficiency and delay.

### v2

@@ D.1 Blockchain Efficiency and Transaction Delay
We define blockchain efficiency as the expected number of blocks in the strongest chain (path) divided by the total number of blocks broadcast, in the limit of many blocks. For viable parameterizations this concentrates at a finite value as the number of block broadcasts becomes large. For purposes of Figs. 5 and 16 we estimate this value once after every block broadcast for blockchains of between $512$ and $1024$ broadcasts (mining events). We discard the values for the first 50% of the chain evolution and average the remaining result. The resulting value is shown as a mean and standard error with respect to 16 independent chains.
@@
For figure 16 we measure 1 pair-delay per chain following each mining event on chains of length from 512 up to 1024. We determine the empirical distribution of $D_{pair}$, discarding statistics from the first half of each chain. We threshold the distribution to determine a transaction delay with 95% confidence, presenting the mean and variance with respect to 16 blockchains as indicated in captions. Although pair-delays measured this way are correlated, for each data point we have at least thousands of effectively independent samples for each data point, so that a 95% quantile is established with reasonable confidence [60].
@@ D.2 Mining Acceleration for Blockchain Analysis
Each mining event requires $O(2^{\mathcal{N}_{\rm zeros}})$ experiments. For purposes of blockchain structural analysis we show methods whereby only $O(1)$ experiments are required per mining event. This allows us to present, with practical compute resources, results on stability for large $\mathcal{N}_{\rm zeros}$, limited primarily by the number of validation experiments rather than the exponentially growing mining timescale.
@@
We first make the approximation in all experiments of perfect network synchronization, discounting the rare event of two miners simultaneously completing a proof-of-work. This approximation is reasonable if the mining rate is small enough.
We employ a hashing method where the overall sign on any hyperplane is uniformly distributed in $\{-1,1\}$. For this reason the probability a nonce describes a broadcastable (valid) block in case (a) is $1/2^{-\mathcal{N}_{\rm zeros}}$; mining is successful if and only if the hyperplane signs $H_{\alpha}=\mathrm{sign}(W_{\alpha})$. We can therefore emulate properly distributed statistics of a fair mining event by first running an experiment to produce $W_{\alpha}$, and then conditionally sampling the signs on hyperplanes to allow mining.
@@
D.3 Bootstrapping Witness Methodology
@@
For a training set of 5 experiments and $1024$ random projections, the distribution of bits (thence validation rates), and distribution of $W_{\alpha}$ (thence confidence in validation) can be established as shown in Figs. 13 and 14. We can fit the witness distributions by Gaussians, and thresholded bit outcomes by Bernouilli distributions. For purposes of accelerating analysis we can then resample these distributions uniformly at random to model a larger set of experiments without use of the QPU experiment each time.
We employ this procedure in Figs. 5 and 16 with results consistent with full blockchain experiments involving finite-number of miners, such as those in Fig. 10. We consider 20480 combinations of model (spin-glass realization) and hyperplanes, for each of the 4 experimental QPUs. Using 32 programmings, we determine the distribution (Gaussian or Bernouilli for witnesses and bits respectively). We can then use these parameterized distributions to generate reasonable witness (and thresholded witness) distributions. For each mining attempt, we uniformly randomly sample a combination of model and hyperplanes - and then for any mining or validation event we choose uniformly at random a QPU and sample from the corresponding distribution.
@@
Qualitative justification for this approach relies on the observation that a spin-glass ensemble operated in our experimental regime is self-averaging. The distributional properties of such witnesses fluctuate little between typical spin-glass disorder realizations. Variation in witnesses values between QPUs (and embeddings) appears well described by Gaussian distributions, and since hyperplanes are nearly orthogonal the correlations in their fluctuations are weak. We plot in Figs. 5 both results obtained by bootstrapping and those obtained by other experiments (either with fixed number of miners, or many miners) with good agreement, validating the approximation. Full simulations (every verification and mining event uses a QPU) are used for all other figures.